OpenStudy (anonymous):

Are these fractions equivalent? 7/10 and 5/6 and 8/12 and 4/6

OpenStudy (across):

Just for fun, let us multiply the top and the bottom of 7/10 by 6: (7*6/10*6)=42/60. Now, let us multiply the top and the bottom of 5/6 by 10: 5*10/6*10=50/60. Now that they have the same denominator, look at their numerators: is 42=50?

OpenStudy (across):

Now do this for the second one.

OpenStudy (anonymous):

Why did you choose 6? I understand it by the fraction tiles better, is tehr a way to explain it that way?

OpenStudy (across):

My goal is to make both fractions have the same denominator. That is why I chose 6 for the first one and 10 for the second one.

OpenStudy (anonymous):

Okay I got it the second one is equilvent.